Cos​​−1x+cos−1y+cos−1z =π and x+y+z =3/2 then prove that x=y=z

Grade:12

1 Answers

Arun
5 years ago
 
Dear student
 
we proceed with reverse side
x = y = z
 
then 3 cos^-1 x = pi
 
cos^-1 x = pi/3
x = ½
 
hence y = z= ½
 
hence 
 
x + y + z = 3/2